OpenStudy (anonymous):

how do you simplify (-27)^-(1/3)

OpenStudy (anonymous):

this means \[\frac{1}{\sqrt[3]{-27}}\]so answer is \[-\frac{1}{3}\]

OpenStudy (anonymous):

the exponent of \[-\frac{1}{3}\] has a minus sign. that means take the reciprocal. the denominator is 3, that means cubed root. numerator is 1 so first power